OpenStudy (anonymous):

A group of five friends spent a total of $18 on a bottle of juice and a package of dried fruit for each of them. Each juice cost $1.25 each. Each package of dried fruit cost the same amount. What was the cost of each package of dried fruit?

OpenStudy (anonymous):

5 * 1.25 = $6.25

OpenStudy (anonymous):

18 - 6.25 = $11.75, which is the amount left to buy 5 dried fruit packages.

OpenStudy (studygurl14):

Use this equation: 1.25(5) + C(5) = 18 where C equals the cost of a package of dried fruit Solve for C

OpenStudy (anonymous):

So 11.75/5 = $2.35 for each package of dried fruit.

OpenStudy (anonymous):

Do you know how to do a tape program for this problem?

OpenStudy (anonymous):

Tape diagram?

OpenStudy (anonymous):

I've never heard of a tape diagram.. Sorry.